# ARise: Towards Knowledge-Augmented Reasoning via Risk-Adaptive Search

### v1

@@ 3.3 Comparison to Learning-based LRMs
Our empirical comparison between base models with ARise and the DeepSeek-R1 distilled models reveals key insights into the effectiveness of test-time search. These learning-based LRMs extract the similar reasoning pattern from DeepSeek-R1. As shown in Figure 3, ARise exhibits a performance advantage over the LRMs, especially on the Qwen model series. While ARise slightly underperforms in comparison to DeepSeek-R1-style reasoning pattern on Llama, it consistently outperforms it on the Qwen models. On average, ARise shows a relative improvement of 4.03%, emphasizing the benefit of our search-based method. These results suggest that while learning-based LRMs like DeepSeek-R1 distilled models provide valuable insights, they have not yet demonstrated the same effectiveness as search-based reasoning.
@@ 3.4 Model Scale
We conducted experiments on the Qwen2.5 series models, spanning a range of parameter scales from 0.5B to 32B, as illustrated in Figure 4. To realistically and prominently reflect the related trends, we selected two more challenging test sets: 2WikiMultihopQA and Musique. We employ Pass@N as the metric to evaluate the upper bound of the success rate, where a problem is considered solved as long as a single surviving node in the tree leads to the correct answer. Pass@1, on the other hand, represents the success rate of the optimal path selected under the guidance of ARise. The results show that Pass@N and the vanilla method exhibit similar trends as model parameters scale, but there is an average of 26.85% significant room for improvement. This indicates that appropriately scaling up inference computation offers substantial potential for enhancing performance. We observed that after the model size reaches 7B parameters, the performance of both the upper bound and naive retrieval tends to saturate, suggesting diminishing returns with further scaling of model parameters. In contrast, ARise demonstrates consistent improvement as model parameters increase. The accuracy of the optimal path selection gradually approaches the upper bound, with the success rate gap between Pass@1 and Pass@N decreasing from 25.00% to 7.25%.
3.5 Ablation Studies
3.5.1 Risk Assessment
Method HotpotQA 2Wiki MusiQue Average
EM F1 EM F1 EM F1 EM F1
MCTS (R-V function) 73.50 75.39 56.50 62.61 40.50 65.98 56.83 67.96
MCTS (vanilla) 71.00 73.58 48.50 60.47 34.50 63.34 51.33 65.80
Finding 5: The Risk-Value function effectively risk-assess reasoning states to guide the tree search process.
3.5.2 Iterations of MCTS
Finding 6: ARise with dynamic risk assessment achieves near-optimal solutions with a relatively low inference cost.
We conduct empirical experiments on the trade-off between inference cost and solution quality. Specifically, we examined the performance of ARise as the number of MCTS iterations increased from 1 to 200. The Figure 5 illustrates the efficiency of ARise in reaching near-optimal solutions with a relatively low inference cost.
Figure 5: Ablation on iterations. We evaluate performance using the average of the EM accuracy and the F1 score.
4 Related works
Test-Time Compute.
Retrieval-Augmented Generation.
5 Conclusion
In this work, we introduce ARise, a novel framework for knowledge-augmented reasoning that addresses the challenges of error propagation and verification bottleneck in complex reasoning tasks. By integrating MCTS with risk-adaptive exploration, ARise enables dynamic and effective reasoning through iterative decomposition, retrieval-then-reasoning steps. Our experiments demonstrate that ARise outperforms a wide range of state-of-the-art methods, and also surpasses the performance of the latest learning-based LRMs equipped with RAG. These results highlight the potential of ARise in advancing open-ended, knowledge-intensive, and complex reasoning tasks in real-world applications.
Limitations
Although our method ARise demonstrates strong performance for knowledge-intensive, and complex reasoning tasks, several limitations remain open for improvement. Our current work focuses test-time search and leaves the training of the policy model for future research. While we have demonstrated the effectiveness of our method in dynamically guiding the solution-search process, a learning-based LRMs could potentially further optimize the trade-off between performance and inference cost. The prompts employed in our experiments were selected based on empirical trials rather than a systematic optimization process. Although these prompts have shown to be effective in our tasks, a more rigorous approach to prompt design and tuning could lead to improved generalization and robustness across diverse scenarios. Our experiments are currently confined to multi-hop question-answering (QA) tasks. The applicability of ARise to other reasoning tasks, such as mathematical problem-solving, code generation, or complex decision-making, remains to be explored. Extending our method to a broader range of reasoning tasks is an important direction for future work.
Appendix A Additional Results
A.1 Discussion on Computational Overhead
We analyze the computational overhead of ARise from two perspectives: (1) its relation to the search space, and (2) its comparison with diverse baselines. All experiments were conducted using two NVIDIA H100 GPUs (80GB), and the cost is primarily measured in terms of reasoning time (in minutes), which reflects real-time efficiency during inference.
Computational Overhead in Relation to Search Space.
(Depth, Width) EM F1 Time (min)
Vanilla / (1,1) 56.00 59.70 10
(3,3) 63.50 65.84 40
(3,4) 65.00 65.69 52
(3,5) 64.50 66.63 66
(3,6) 65.00 64.91 83
(4,4) 66.50 65.36 119
(4,5) 66.00 67.01 168
(4,6) 67.00 66.29 266
(5,4) 63.50 65.52 189
(5,5) 66.00 66.24 413
(5,6) 65.50 65.63 770
@@ A.1 Discussion on Computational Overhead
Table 3: Computational overhead in relation to the search space. As the search space expands in depth and width, performance improves but with diminishing returns and significantly increased reasoning time. The (4,5) setting achieves a good trade-off and is adopted in our main experiments.
@@
We conducted experiments on the HotpotQA dataset using the Llama3.1-8B-Instruct model. The results are under different search space configurations—defined by search depth and width. As shown in the Table 3, expanding the search space leads to performance improvements (in EM and F1 scores), especially when increasing from a vanilla setting (depth 1, width 1) to moderate configurations. However, these improvements diminish as the depth and width increase further, while the computational overhead escalates rapidly. For instance, increasing the depth from 3 to 4 and the width from 4 to 6 results in a reasoning time increase from 52 to 266 minutes, with only marginal performance gain. Based on this observation, we adopted a configuration of depth = 4 and width = 5 for our main experiments, which strikes a practical balance between accuracy and computational cost. It is worth noting that the optimal search space configuration may vary across instances. Different questions demand varying depths of reasoning and degrees of knowledge integration, often corresponding to different numbers of reasoning hops. Thus, adaptively determining the search space based on task complexity remains an open research problem. Even state-of-the-art reasoning models still face the challenge of choosing between wider and deeper reasoning paths. We discuss this further in the Limitations and identify it as a promising direction for future work.
Computational Overhead Comparison with Baselines.
@@
A.2 Impact of Retrieval System Quality
@@
Table 5: Impact of Retriever Type. Dense retriever (BGE) outperforms sparse retriever (BM25) across all datasets, with more notable gains on challenging benchmarks.
@@ B.4 Further Details for Prompts
Algorithm 1 ARise
Require: original problem $q$; prompt $p_{d}$ for decomposition; prompt $p_{r}$ for reasoning
Require: LLM $f(\cdot)$; external knowledge retrieval model $g(\cdot)$
Require: selecte $Se(\cdot)$; expand $Ex(\cdot)$; simulate $Si(\cdot)$; back-propagate $Ba(\cdot)$
Require: prompt $p_{g}$ for problem generation; value function $V(\cdot)$; get best path $P(\cdot)$
   $\mathbf{s_{0}};\mathbf{t}\leftarrow q$triangleright Generate root node and tree
   for $i=1,\dots,T$ do $\triangleright$ Perform T MCTS iterations
    $\mathbf{s_{i}}\leftarrow Se(\mathbf{s_{0}},\mathbf{t})$triangleright Select the best node
    $d_{i}\leftarrow f(\mathbf{s_{i}},q,p_{d})$triangleright Decompose the problem
    $z_{i}\leftarrow g(d_{i})$triangleright Retrieve relevant documents
    $r_{i}\leftarrow f(z_{i},d_{i},p_{r})$triangleright Reason forward
    $\mathbf{s_{i+1}}\leftarrow\mathbf{s_{i}},r_{i},V(f(r_{i},q,p_{g}))$triangleright Generate new nodes
    $\mathbf{t}\leftarrow Ex(\mathbf{s_{i+1}})$triangleright Append nodes to the tree
    $\mathbf{t}\leftarrow Si(\mathbf{s_{i}})$; $Ba(\mathbf{s_{i}},\mathbf{s_{0}})$triangleright Simulate and Back-propagate
   end for
   return $P(\mathbf{s_{0}},\mathbf{t})$triangleright Get best path

### v2

@@ 3.3 Comparison to Learning-based LRMs
Our empirical comparison between base models with ARise and the DeepSeek-R1 distilled models reveals key insights into the effectiveness of test-time search. These learning-based LRMs extract similar reasoning patterns from DeepSeek-R1. As shown in Figure 3, ARise exhibits a performance advantage over the LRMs, especially on the Qwen model series. While ARise slightly underperforms in comparison to DeepSeek-R1-style reasoning pattern on Llama, it consistently outperforms it on the Qwen models. On average, ARise shows a relative improvement of 4.03%, emphasizing the benefit of our search-based method. These results suggest that while learning-based LRMs like DeepSeek-R1 distilled models provide valuable insights, they have not yet demonstrated the same effectiveness as search-based reasoning.
@@ 3.4 Model Scale
We conducted experiments on the Qwen2.5 series models, spanning a range of parameter scales from 0.5B to 32B, as illustrated in Figure 4. To realistically and prominently reflect the related trends, we selected two more challenging test sets: 2WikiMultihopQA and Musique. We employ Pass@N as the metric to evaluate the upper bound of the success rate, where a problem is considered solved as long as a single surviving node in the tree leads to the correct answer. Pass@1, on the other hand, represents the success rate of the optimal path selected under the guidance of ARise. The results show that Pass@N and the vanilla method exhibit similar trends as model parameters scale, but there is an average of 26.85% significant room for improvement. This indicates that appropriately scaling up inference computation offers substantial potential for enhancing performance. We observed that after the model size reaches 7B parameters, the performance of both the upper bound and naive retrieval tends to saturate, suggesting diminishing returns with further scaling of model parameters. In contrast, ARise demonstrates consistent improvement as model scale increases. The accuracy of the optimal path selection gradually approaches the upper bound, with the success rate gap between Pass@1 and Pass@N decreasing from 25.00% to 7.25%.
3.5 Computational Overhead
The computational overhead is measured in terms of reasoning time (in minutes).
3.5.1 Overhead in Relation to Search Space.
Finding 5: Moderate search space expansion boosts performance over minimal settings, but returns diminish rapidly as overhead surge.
@@ 3.5 Computational Overhead
Figure 5: Computational overhead in relation to the search space. As the search space expands in depth and width, performance improves but with diminishing returns and significantly increased reasoning time. The (4,5) setting achieves a good trade-off and is adopted in our main experiments.
@@
We conducted experiments on the HotpotQA dataset using the Llama3.1-8B-Instruct model. The results are under different search space configurations—defined by search depth and width. As shown in Figure 5, expanding the search space leads to performance improvements (in EM and F1 scores), especially when increasing from a vanilla setting (depth 1, width 1) to moderate configurations. However, these improvements diminish as the depth and width increase further, while the computational overhead escalates rapidly. For instance, increasing the depth from 3 to 4 and the width from 4 to 6 results in a reasoning time increase from 52 to 266 minutes, with only marginal performance gain. Based on this observation, we adopted a configuration of depth = 4 and width = 5 for our main experiments, which strikes a practical balance between accuracy and computational cost. It is worth noting that the optimal search space configuration may vary across instances. Different questions demand varying depths of reasoning and degrees of knowledge integration, often corresponding to different numbers of reasoning hops. Thus, adaptively determining the search space based on task complexity remains an open research problem. Even state-of-the-art reasoning models still face the challenge of choosing between wider and deeper reasoning paths. We discuss this further in Limitations and identify it as a promising direction for future work.
3.5.2 Overhead Comparison with Baselines.
@@
Finding 6: ARise achieves the best performance but incurs higher computational overhead due to its search-based paradigm.
@@
3.6 Ablation Studies
3.6.1 Risk Assessment
Method HotpotQA 2Wiki MusiQue Average
EM F1 EM F1 EM F1 EM F1
MCTS (R-V function) 73.50 75.39 56.50 62.61 40.50 65.98 56.83 67.96
MCTS (vanilla) 71.00 73.58 48.50 60.47 34.50 63.34 51.33 65.80
Finding 7: The Risk-Value function effectively risk-assess reasoning states to guide the tree search process.
3.6.2 Iterations of MCTS
Finding 8: ARise with dynamic risk assessment achieves near-optimal solutions with a relatively low inference cost.
We conduct empirical experiments on the trade-off between inference cost and solution quality. Specifically, we examined the performance of ARise as the number of MCTS iterations increased from 1 to 200. Figure 6 illustrates the efficiency of ARise in reaching near-optimal solutions with a relatively low inference cost.
Figure 6: Ablation on iterations. We evaluate performance using the average of EM accuracy and F1 score.
4 Related Works
Test-Time Compute.
Retrieval-Augmented Generation.
5 Conclusion
In this work, we introduce ARise, a novel framework that addresses the challenges of error propagation and verification bottleneck in open-ended, knowledge-intensive, and complex reasoning tasks. By integrating Monte Carlo Tree Search with risk-adaptive exploration, ARise enables dynamic and effective reasoning through iterative problem-decomposition and retrieval-then-reasoning steps. Our experiments demonstrate that ARise outperforms a wide range of state-of-the-art methods, and also surpasses the performance of the latest learning-based large reasoning models (LRMs) equipped with RAG. These results highlight the strong potential of ARise in advancing open-ended, knowledge-intensive, and complex reasoning tasks across various real-world applications.
Limitations
Although our method ARise demonstrates strong performance for knowledge-intensive, and complex reasoning tasks, several limitations remain open for improvement. Our experiments are currently confined to multi-hop question-answering (QA) tasks. The applicability of ARise to other reasoning tasks, such as mathematical problem-solving, code generation, or complex decision-making, remains to be explored. Extending our method to a broader range of reasoning tasks is an important direction for future work. There is also a need for more systematically designed prompts to ensure generalization and robustness across diverse scenarios. Moreover, existing search-based paradigms mainly rely on post-trained reward models for verification. These models are trained on specially annotated data to learn how to score reasoning states. While this approach has shown progress in mathematics (a relatively closed domain) Lightman et al. (2024); Luo et al. (2024); Wang et al. (2024), static, post-trained reward models struggle with open-ended, knowledge-intensive tasks, where external knowledge is dynamically involved. Such models cannot accurately assess new knowledge unseen during post-training, limiting their generalization. We argue that a generalizable design of rewards is crucial for future developments. Our work presents an initial attempt in this direction, and we will further explore this line of research. Last but not least, the search space for specific questions is predefined, lacking the ability to adapt efficiently to varying reasoning complexity or knowledge density. This points to an open challenge in the study of efficient reasoning: how to achieve an effective trade-off between broad and deep search, and how to reduce redundancy in the reasoning process. We leave this as an avenue for future work.
Acknowledgement
We thank all the anonymous reviewers and area chair for their valuable feedback throughout the review process. This work is supported by Shanghai Artificial Intelligence Laboratory.
Appendix A Impact of Retrieval System Quality
@@
Table 4: Impact of Retriever Type. The dense retriever (BGE) outperforms the sparse retriever (BM25) across all datasets, with more notable gains on challenging benchmarks.
@@ B.4 Further Details for Prompts
Problem Decomposition Your task is to decompose the original question into one smaller sub-question based on the Intermediate answer and Observation. The decomposed process is encouraged to be done from multiple perspectives. Output a thought to reason the original question, and output one sub-question that you think is appropriate to solve next. DO NOT REPEAT the question and DO NOT try to answer the question. The output format is limited to: Thought: … Sub-question: … Here, the "…" indicates omitted output information that you need to fill in. Original question: {original question} Intermediate answers: {reasoning state} Observation: {retrieved documents} Output:
Intermediate Answer Generation Your task is to answer the following question using provided supporting facts. The output answer should be a complete declarative sentence, rather than directly outputting phrases or words. DO NOT use pronouns in the sentence. Specially, if no provided supporting facts, just output "No directly relevant facts found." and nothing else. Question: {sub-question} Supporting facts: {retrieved documents} Output:
Final Answer Generation Your task is to answer the original question based on the intermediate answers. Output the final answer directly and nothing else. Original question: {original question} Intermediate answers: {reasoning state} Output:
Risk Assessment Given intermediate answer containing the facts about the original question, which is unknown, your task is to infer what the original question might have been. Output the most likely original question directly and nothing else. Example 1: Intermediate answer: Muhammad Ali was 74 years old when he died. Alan Turing was 41 years old when he died. The original question might be: Who lived longer, Muhammad Ali or Alan Turing? Example 2: Intermediate answer: Craigslist was founded by Craig Newmark. Craig Newmark was born on December 6, 1952. The original question might be: When was the founder of craigslist born? Intermediate answers: {reasoning state} The original question might be:
LLM-as-Verifier Given a question, your task is to determine the consistency score of its decomposed sub-questions and corresponding intermediate answers with the original question. Directly output JUST A NUMBER between 0 and 10 to represent the consistency score. DO NOT output anything else. Original question: {original question} Sub-questions: {sub-questions} Intermediate answers: {reasoning state} Output:
ARise Simple Case STATE 1: depth: 1; action: decompose original question: When was the last time Peter Till’s team beat winner of 1894-95 FA Cup in SC? sub-questions: Who was the winner of FA Cup? intermediate-answer: reasoning state: STATE 2: depth: 1; action: retrieve-then-reason original question: When was the last time Peter Till’s team beat winner of 1894-95 FA Cup in SC? sub-questions: Who was the winner of FA Cup? – retrieve: {supporting facts} intermediate-answer: The winner of the 1894- 95 FA Cup is Aston Villa. reasoning state: The winner of the 1894- 95 FA Cup is Aston Villa. STATE 3: depth: 2; action: decompose original question: When was the last time Peter Till’s team beat winner of 1894-95 FA Cup in SC? sub-questions: What was the sports team of Peter Till? intermediate-answer: reasoning state: The winner of the 1894- 95 FA Cup is Aston Villa. STATE 4: depth: 2; action: retrieve-then-reason original question: When was the last time Peter Till’s team beat winner of 1894-95 FA Cup in SC? sub-questions: What was the sports team of Peter Till? – retrieve: {supporting facts} intermediate-answer: Peter Till’s sports team is Birming-ham City. reasoning state: The winner of the 1894- 95 FA Cup is Aston Villa. Peter Till’s sports team is Birming-ham City. STATE 5: depth: 3; action: decompose original question: When was the last time Peter Till’s team beat winner of 1894-95 FA Cup in SC? sub-questions: When was the last time Birming-ham City beat Aston Villa in SC? intermediate-answer: reasoning state: The winner of the 1894- 95 FA Cup is Aston Villa. Peter Till’s sports team is Birming-ham City. STATE 6: depth: 3; action: retrieve-then-reason original question: When was the last time Peter Till’s team beat winner of 1894-95 FA Cup in SC? sub-questions: When was the last time Birming-ham City beat Aston Villa in SC? – retrieve: {supporting facts} intermediate-answer: The last time Birming-ham City beat Aston Villa was in March 2005. reasoning state: The winner of the 1894- 95 FA Cup is Aston Villa. Peter Till’s sports team is Birming-ham City. The last time Birming-ham City beat Aston Villa was in March 2005. STATE 7: depth: 4; action: answer original question: When was the last time Peter Till’s team beat winner of 1894-95 FA Cup in SC? sub-questions: intermediate-answer: reasoning state: The winner of the 1894- 95 FA Cup is Aston Villa. Peter Till’s sports team is Birming-ham City. The last time Birming-ham City beat Aston Villa was in March 2005. final answer: March 2005 NOTE: Decomposition is based on original question, reasoning state, and retrieved documents.